OpenStudy (anonymous):

Which answer shows this equation in standard form? 7 – 3(x – y) = –5x + 2 A. –8x + 3y = –5 B. –2x – 3y = 5 C. 2x + 3y = –5 D. 2x – 3y = –5

OpenStudy (anonymous):

@leonardo0430 @dan815 @Compassionate

OpenStudy (anonymous):

@robtobey @campbell_st @e.mccormick

OpenStudy (compassionate):

7 – 3(x – y) = –5x + 2 Standard form is Ax +By = C So we need to solve first 7 – 3(x – y) = –5x + 2 Distribute the -3 first. Then subtract 7 from both sides.

OpenStudy (anonymous):

soo -3-7 – 3(x – y) = –5x + 2-3

OpenStudy (anonymous):

then subtract 7 so C

OpenStudy (compassionate):

7 – 3(x – y) = –5x + 2 7 - 3x + 3y = -5x + 2 -3x + 3y = -5x - 5 Add five x. 2x + 3y = -5 Yup!

OpenStudy (anonymous):

@leonardo0430 you were right thx @Compassionate thank you can you still help me with 3 more questions

OpenStudy (anonymous):

What is the x-intercept of the line with the equation 1/3x+y=-15 ?

OpenStudy (anonymous):

@sammixboo @dan815 @panda_lover

OpenStudy (compassionate):

To find the x-intercept, solve for x.

OpenStudy (compassionate):

Or, we can do it graphically, which is easier. So take (1/3)x + y = -15 To find the x-intercept, replace y with zero. (1/3)x = -15 Multiply by 1/3 3/1 * (1/3)x = 3/1 * -15 Now what will we do?

OpenStudy (compassionate):

Generally, if you want to find the x-intercept, replace y with 0 and solve for x. Vice-versa, for the y-intercept, replace x with 0 and solve for y.

OpenStudy (anonymous):

is it 5,0 im horrible a this subject

OpenStudy (compassionate):

No no. \[\frac{ 3 }{ 1 } \times \frac{ 1 }{ 3 }x = -15 \times \frac{ 3 }{ 1 }\] On the left, we just multiply straight across. 3 x 1 and 1 x 3 will equal 3/3, which is just 1. On the right, we have -15 times 3/1. This can be expressed as: \[\frac{ -15 }{ 1} \times \frac{ 3 }{ 1}\] Because putting a number over 1 does not change its value. So, when you see something like: \[4 \times \frac{ 3 }{ 2 }\] What its really saying is: Four times Three, divided by 2. The same concept applies here. So try again.

OpenStudy (anonymous):

so we want to -15x3/1

OpenStudy (compassionate):

Yes! \[\frac{ 1 }{3 }x + y = -15 \]\[\frac{ 1 }{ 3 }x + 0 = -15\] [ Since we are solving for x, replace y with 0. Thus we are left with] \[\frac{ 1 }{ 3 }x = -15\][Now, the final step is solving for x. Since 1/3 is being multiplied by x, to do the opposite, we want to multiply by the inverse of 1/3, which is 3/1. Multiplying by the inverse of a number is the same as dividing it.] Can you do it from here? Remember, multiply both sides by 3/1

OpenStudy (anonymous):

but i get -45 when i do -15x3, and i dont think its right

OpenStudy (anonymous):

wait i think i got it is it 5,0

OpenStudy (anonymous):

@Compassionate

OpenStudy (compassionate):

\[-15 \times \frac{ 3 }{ 1} = \frac{ -15 \times 3 }{ 1 }\]

OpenStudy (anonymous):

-15 x 3=-45 -45/1=45

OpenStudy (compassionate):

Well, it would be negative 45.

OpenStudy (anonymous):

so is it 45,0 i do understand you i just cant plug it in

OpenStudy (compassionate):

No, no. You don't need to plug in anything. Just replace y with 0, then solve for x. We just discovered x = -45 -45/1 = -45. Not 45. So the points are (-45, 0)

OpenStudy (anonymous):

so i was right thats what i thought but then got confused sorry for being so annoying thx for the help and for being patient with me

OpenStudy (compassionate):

Mhm. Anything else you need help with?

OpenStudy (anonymous):

umm..... let me see

OpenStudy (anonymous):

yah i need help on equations and graphs

OpenStudy (compassionate):

Shoot.

OpenStudy (anonymous):

but thx ill open another page

OpenStudy (leonardo0430):

no problem :) sorry i didnt have time to explain :l

OpenStudy (anonymous):

its ok thx